OpenStudy (anonymous):

The ratio of two numbers is 7 to 5. The sum of the numbers is 24. What are the numbers?

OpenStudy (mertsj):

\[\frac{x}{y}=\frac{7}{5}\] \[x+y=24\]

OpenStudy (mertsj):

Solve the system